### v2

# Modelling Realistic Multi-layer devices for superconducting quantum electronic circuits

Abstract
In this work, we present a numerical model specifically designed for 3D multilayer devices, with a focus on nanobridge junctions and coplanar waveguides. Unlike existing numerical models, ours does not approximate the physical layout or limit the number of constituent materials, providing a more accurate and flexible design tool. We calculate critical currents, current phase relationships, and the energy gap where relevant. We validate our model by comparing it with published data. Through our analysis, we found that using multilayer films significantly enhances control over these quantities. For nanobridge junctions in particular, multilayer structures improve qubit anharmonicity compared to monolayer junctions, offering a substantial advantage for qubit performance. For coated multilayer microwave circuits it allows for better studies of the proximity effect, including their effective kinetic inductance.
Most superconducting quantum technologies rely on Josephson Junctions RevLuk79; RevGol2004; intrtrasmon and coplanar waveguides (CPW) CPW1; wallraffcpw for their operation. To meet the rigorous standards required for such applications, the simulation of these components becomes an essential part of the fabrication process. This work develops a numerical model tailored for coplanar waveguides (CPW) CPW1; wallraffcpw and nanobridge junctions, which offer a promising alternative to traditional tunnel junctionsdevoret2004superconducting; hertzberg2021laser in superconducting quantum devices, such as qubits Arute2019; Miyajima:18 and Single Flux Quantum (SFQ) tolpygo2018superconductor due to several key advantages. Unlike tunnel junctions, which rely on an oxide layer that can degrade performance by creating localized states, energy loss, and reduced coherence Burnett2019; PhysRevLettMartin, nanobridges use a small metallic bridge to connect electrodes directly. This structure removes the oxide layer entirely, minimizing unwanted capacitance that interferes with precise qubit engineering and eliminating fabrication complexities that often hinder integration with other on-chip components collins. These advantages make nanobridges particularly suitable for superconducting qubits. Previous theoretical works were limited to 3-D simulations3dFEMSNS that did not include applications in quantum technologies or 2-D nanobridge designs Vijayjunction; Ragazzidikupr_2021; MiniaturizationofJosephsonJunctions; 2023_Kup_Gol which did not consider more elaborate shapes such as rounded edges, which have an effect on the junction properties as it will be shown later in the paper. Therefore, we aim to overcome these limitations and interface our model with the standard design process of superconducting quantum devices. We also explore additional scenarios where our model can be applied, such as calculating the effect of the proximity effect in interfaces between superconductors and other metals, to establish the reduction of the energy gap in the structure which can have a detrimental effect on qubits’ coherence time. This area has attracted enormous interest, especially since recent studies have shown that surface encapsulation techniques, involving passivating the surface of superconducting materials like niobium to prevent the formation of lossy oxides, can significantly enhance qubits’ performance encapsulation. This broader applicability underscores the versatility of our model in enhancing the design and performance of sophisticated superconducting quantum devices, paving the way for more reliable and scalable quantum technologies.
A nanobridge Josephson junction consists of two superconducting electrodes connected by a small constriction, usually of order of a few tens of nanometers. These constrictions may consist of the same material or a different one, with the latter referred to as SS’S or SNS junctions, where S represents a superconductor and S’ or N represents the filling material that connects the two electrodes, which may be another superconductor (S’) or a normal metal (N).
Under the assumption that the magnetic effects are negligible and all materials satisfy the dirty limit condition (), where is the mean free path, and is the superconducting coherence length, we can describe the current transport properties of this type of junction with the Usadel equations RevGol2004; Usadel70; Kurpianov1988; Ragazzidikupr_2021.
Here, is the matrix Usadel Green’s function in layer . The term represents the energy, and is the diffusion coefficient in layer , where is the Fermi velocity. The term is the third Pauli matrix, and is the pair potential matrix, defined as
The coherence length, the diffusion coefficients and the critical temperature () of a superconductor are related by
To determine stationary properties of the junction, such as current densities, the Usadel equations are solved for imaginary energies , where the Matsubara frequencies are given by . In this regime, it is convenient to parametrize the Green’s function using the -parametrization:
which must be solved for and using the self-consistent equation:
and the boundary conditions between different materials:
where is the normal at the boundary, represents the critical temperature of the superconducting electrodes, is the Boltzmann constant and . Defining as the resistivity of material , and represent the suppression parameter and the boundary resistance parameter, respectively. The suppression parameter describes how much the superconductivity is reduced going from S to the weak link (no suppression means ), while the boundary resistance parameter describes the effect of the surface resistance () at the boundary between and . Far from the bridge on the surfaces not connected to the layer, this boundary condition is applied:
where is the Bardeen–Cooper–Schrieffer (BCS) energy gap and is the phase difference between the bulk electrodes Ragazzidikupr_2021. The superconducting current is calculated as follows:
where is the resistivity of the whole junction in the normal state and is the cross-sectional area.
We will, now, discuss how we reformulated and solved the set of PDE in Eqs.(4, 5). The set of PDE is not only non-linearly dependent on the parameter but also non-differentiable in , because of the term . This added difficulty forced us to separate the real and imaginary parts of the equations to unsure convergence with the Newton method, actually creating a set of coupled equations that in the weak form become:
and for the imaginary part:
Here, and are the test function sassociated respectively to and FEM, represents the surface separating the materials and . The equations for the second material, denoted as , are identical except for differences in the integrals over due to distinct boundary conditions as outlined in Eq.(6). The equations were solved in sfepy sfepy, by creating new customs non-linear functions.
[…]
Superconducting quantum computing circuits require Josephson junctions with small critical currents, of the order of 10 to 100 nA, and CPRs as close as possible to a sinusoid in order to obtain non-linear energy levels with high anharmonicitynanobridgequbits. For SFQs the constrain of low critical currents is lifted but the shape of the CPR still plays a crucial role in the generation of the pulses RSFQ. Moreover, to stay within the Josephson effect, the dimensions of the bridge should be less than 3 times the coherence length RevGol2004. For these reasons, different layouts have been numerically investigated as shown in Fig. 1, also taking into account realistically rounded edges that can come for an eventual etching process collins, the lateral dimension W of the nano-bridge was chosen to be and for all the simulated devices. We calculated the current-phase relationship (CPR) for the following use-case scenarios:
Planar junctions with rectangular and rounded cross-sections
Variable Thickness Nanobridge SS’S Junctions (VTB) with rectangular and rounded cross-sections
These devices are illustrated in Fig.1. The simulation results are presented in Fig.2, where we demonstrate how the rounded edges reduce the critical current in the junction, even after normalizing for the reduced cross-section. Additionally, in the case of the planar junction, the rounded edges slightly shift the maximum of the CPR to the left. This occurs because a reduced cross-section weakens the connection between the electrodes while maintaining the ratio, bringing the junction closer to the ideal case described by the one-dimensional Usadel equations, known as the KO-1 model KUPRIYANOV1981, which is also shown in the figure. Compared to the KO-1 model, the CPR exhibits lower values and increased skewness, which can be attributed to the reduction of the energy gap in the bridge area relative to the bulk electrodes, as illustrated in the inset. This reduction does not occur in the KO-1 model, where only the bridge is simulated with the assumption that the energy gap at its boundaries matches the bulk value. However, VTB junctions, for the same set of parameters, produce less skewed CPRs, which can improve qubit anharmonicity as discussed previously. This improvement is primarily due to the variable thickness design, where the electrodes act as a robust phase reservoir, leading to a more significant phase change across the bridge rather than within the electrodes, compared to planar junctions. This result aligns with the findings in Ref. Vijayjunction. The use of different materials enables better control over the device’s critical current by selecting a bottom material with the appropriate resistivity. We further validated our model by comparing it with the experimental work in Ref. VijaySquid, where aluminium VTB and planar nanobridges of different lengths were measured, as shown in Fig. 2 (b). To do this, we first estimated the resistivity from the IV curve of a 75 nm long VTB nanobridge, and then plotted the simulated data obtained by using both the reported coherence length of 40 nm and an adjusted value of 110 nm, found in Ref. propertiesAl, which aligned more closely to the experimental data. This adjustment is based on the assumption that the coherence length might be larger than the reported one, calculated by using the film’s resistivity that could have been made at room temperature, thus lowering the value of . This comparison underscores the accuracy of our model and validates our design choices.
We are now using our numerical model to optimize the calculation of energy levels in superconducting qubits. The Hamiltonian for the transmon qubit with a junction featuring a generic current-phase relation (CPR) is given by:
while the Hamiltonian for the fluxonium qubit is fluxoniumtrue; Fluxonium; FluxoniumwithEjvalue:
In the case of the fluxonium, the following constraints need to be imposed to differentiate it from other types of inductively shunted devices FluxoniumwithEjvalue, and , which means that the critical current of the nanobridge junction in these devices should be limited to less than 50 nA. Although this is a small value for a nanobridge junction, it can be achieved by utilizing a highly resistive layer. From the CPR in Fig.2 we calculated the associated Josephson energy as nanobridgequbits:
To solve the equation we incorporated in scqubits Scqubits and qiskit metal qiskit a custom Hamiltonian in the phase basis and solved the eigenvalue problem with scipy.linalg.eigh SciPy following the approach in Ref. nanobridgequbits. The eigenenergies of the Hamiltonian are solved for different realistic ratios of and then used to calculate the anharmonicities plotted in Fig.3. As expected the anharmonicity reduces going from a tunnel junction (sinusoidal CPR) to the CPR we obtained with our model; this aligns with previous results nanobridgequbits; zlatkoCPRqubit. However, the low levels of planar Junctions’ anharmonicity are recovered using a multilayer design thanks to its less skewed CPR.
| Material | thickness [nm] | cm] | [nm] | [K] | [] | References | ||
| Au | 10 | 20 | 88 | 0.16 | 0 | 95 | 0.97 | Refs. meanfreepathGold; fewmetalselectronpath |
| TiN | 15 | 300 | 53.6 | 0.02 | 4.5 | 99 | 0.99 | Refs. TiNnanobridge; TiNproperties; ALDTiN; propertiesTiN |
| Al | 10 | 1.5 | 110 | 1.69 | 1.2 | 97 | 0.69 | Ref. propertiesAl |
| Ta | 10 | 21 | 77 | 0.17 | 4.4 | 99 | 0.96 | Refs. propertiesTa; Tafermivel |
The same equation can be employed in any situation where a superconductor and another metal are in contact. Recently, a method known as encapsulation has been utilized in Ref. encapsulation to enhance qubits’ lifetime by depositing a normal metal or another superconductor over the superconducting material, in order to mitigate the detrimental effects of surface oxides. However, due to the proximity effect, the energy gap () decreases in the stronger superconductor and increases in the weaker one causing quasiparticles to start forming at lower drive frequencies. Since it is necessary to satisfy the condition throughout the device tinkham2004introduction, they are more susceptible to environmental noise in a cryostat noisecryo. This makes it crucial to carefully choose the encapsulating layer for superconducting devices, taking into account the material’s thickness and resistivity, as described by the relation . In Fig. 4, we analyze the proximity effect in a superconducting CPW capped with a normal metal, observing that an increase in leads to a decrease in the superconducting gap, consistent with expectations. In Tab. 1 and Fig. 5, we present calculations of the reduction in for the materials listed in the referenced paper. This reduction in leads to an increase in the density of quasiparticles, which in turn reduces the internal quality factor of the resonator and, for qubits, shortens the relaxation time Decoqubit. Therefore, to minimize quasiparticle formation and associated losses, it is advisable to select materials with a high critical temperature () and/or values of resistivity and coherence length such that <1, such as tantalum and TiN. Furthermore, the energy gap was also used to calculate the kinetic inductance of the film integrating in parallel the following equationAnnunziata; Kineticmodel:
The ratio of kinetic inductance with and without the capping layer is reported in Table 1. The materials that exhibit the most significant changes in kinetic inductance are those showing the largest variations in their energy gap, as well as aluminium due to its low resistivity, as expected. Our findings indicate that even when the variation in the energy gap is not substantial for the listed materials, resistivity still plays a major role in determining the total film’s kinetic inductance, even for small capping-layer thicknesses.
In conclusion, we have developed a detailed numerical model that addresses the nanobridge Josephson junction, coplanar waveguide (CPW) resonator, and qubit as integrated elements in superconducting quantum circuits. Our model overcomes the limitations of earlier approaches, which often focused on simplified geometries or treated these components in isolation. By incorporating realistic 3D geometries, rounded edges, and material considerations, we provide a simulation framework that closely mirrors the actual devices used in experiments. The current-phase relationships (CPRs) derived from our simulations demonstrate how the critical current and energy gap are influenced by the nanobridge’s shape and material composition, providing valuable insights into optimizing qubit designs for improved anharmonicity and coherence times. We validated our model by comparing it with experimental data, showing a strong agreement. The flexibility of our model allows it to be applied not only to standard Josephson junctions but also to more complex scenarios like proximity-effect-driven reductions in energy gaps at material interfaces in superconducting CPWs, which are crucial for qubit performance. Overall, our work provides a robust tool for the design and optimization of superconducting qubit circuits. This advancement paves the way for the development of more reliable and scalable quantum devices. Future work will focus on the experimental validation of the devices presented in these designs.
Supplementary Material
This section presents additional simulations exploring different combinations of materials forming the nanobridge junction, along with an example demonstrating the convergence of the simulation as the number of mesh elements increases in the 3D model.
ACKNOWLEDGEMENTS
This work was supported by the UK Engineering and Physical Sciences Research Council (EPSRC) grant EPIT025746/1 for the University of Glasgow. The authors acknowledge V. Georgiev, L. Kaczmarczyk and A. Shvarts for the fruitful discussions.
AUTHOR DECLARATIONS
CONFLICT OF INTEREST
The authors have no conflicts of interest to disclose.
AUTHOR CONTRIBUTIONS
Giuseppe Colletta: Conceptualization (equal), Software (lead), Methodology (lead), Writing/Original Draft Preparation (lead), Writing – review and editing (equal). Susan Johny: Conceptualization (equal), Writing – review and editing (equal). Jonathan A. Collins: Conceptualization (equal), Writing – review and editing (equal). Alessandro Casaburi: Conceptualization (equal), Writing – review and editing (equal), Funding Acquisition (equal). Martin Weides: Conceptualization (equal), Writing – review and editing (equal), Funding Acquisition (equal).
DATA AVAILABILITY
The data that support the findings of this study are available from the corresponding author upon reasonable request.